calculate the pH of 0.32 M Naoh

Grade:

1 Answers

Gaurav
9 years ago
pOH = -log[0.32] = - log[3.2\times 10^-1]
= – log 3.2 +1 = -0.505 +1 = 0.5
pH = 14- pOH = 14-0.5 = = 13.5